Given that d/dx(xlnx−x)=lnx, find the volume of the solid obtained by rotating, about the x-axis, the graph of the function y=√(lnx) from the x-intercept and x=e^2.
1
Expert's answer
2020-05-25T20:47:52-0400

Since "\frac{d}{dx}(xlnx-x)=lnx," then "\int lnx dx= xlnx-x+C"

Volume of the solid obtained by rotating can be found by formula: "\pi \int y^2dx"



Boundaries of integration are "x=1" and "x=e^2".

So:

"V=\pi \int\limits_1^{e^2} (\sqrt{lnx})^2dx=\pi\int\limits_1^{e^2}lnxdx=\pi (xlnx-x)\bigg|_1^{e^2}=\\
\pi(e^2lne^2-e^2)-\pi(1ln1-1)=\pi(e^2+1)"